OpenStudy (anonymous):

sin x (csc x-sin x) = cos^2 x

OpenStudy (anonymous):

Do you know the trigonometric functions?

OpenStudy (anonymous):

i dont understand this at all. i cant do math

OpenStudy (anonymous):

Ok Ill show you some trigonometric functions to make this breeze.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

give me one sec to gather

OpenStudy (anonymous):

thats fine

OpenStudy (anonymous):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html check out this site, it will give you all the info you need.

OpenStudy (anonymous):

You question asks you to solve for sin x (csc x-sin x) = cos^2 x Now lets work out all that is in parenthesis

OpenStudy (anonymous):

If you follow the website I gave you, you would know that cscx= \[\frac{ 1 }{ sinx }\] and sinx in return is sinx= \[\frac{ 1 }{cscx }\] So Now we have \[\frac{ sinx }{ 1 }(\frac{ 1 }{ sinx}-\frac{ 1 }{ cscx })=\cos^2x\]

OpenStudy (anonymous):

I made an error. instead of \[\frac{ sinx }{ 1 }\] Its 1/cscx Now distribute that to everything in parenthesis. You get sinx^2-1=cosx^2 subtract cosx^2 on both sides. Then add 1 on both sides. Your final answer is \[sinx^2-cosx^2=1\] which is an identity.

OpenStudy (anonymous):

you are a life saver. thank you!

OpenStudy (anonymous):

No Problem!